OpenStudy (anonymous):

Factor out Monomial Factors x^-1 -2+x= x^-1(

OpenStudy (anonymous):

Hey it's still me. Do you have any ideas?

OpenStudy (dan815):



OpenStudy (anonymous):

What is it?

OpenStudy (jhannybean):

you have \(\large x^{-1}-2-x\) Lets rewrite it with positive powers. \[\large \frac{1}{x}-2-x\] Make everything a common fraction, as in, set all the bases to x. \[\large \frac{1}{x}-2\cdot \frac{x}{x} - x\cdot \frac{x}{x}\] simplify and rewrite \[\large \frac{1}{x} - \frac{2x}{x}-\frac{x^2}{x}\] make it all one fraction \[\large\frac{1-2x-x^2}{x}\] Can you finish this up? :)

OpenStudy (anonymous):

That's great thank you!

OpenStudy (jhannybean):

No problem! Good luck :)